a chain of fast food restaurants wants to compare the average service times at four different restaurants. They want to conduct a hypothesis test to determine if all four means are the same or not. If n = 10 observations are taken at each of the four restaurants, then the degrees of freedom are df1 = 39 and df2 = 3.
Answer #1

No, this statement is False.

df1 = Number of groups - 1 = 4 - 1 = 3

df2 = Total sample - Number of groups = 4*10 - 4 = 36